# Model Synthesis for Zero-Shot Model Attribution

### v2

@@ Abstract
Nowadays, generative models are shaping various fields such as art, design, and human-computer interaction, yet accompanied by challenges related to copyright infringement and content management. In response, existing research seeks to identify the unique fingerprints on the images they generate, which can be leveraged to attribute the generated images to their source models. Existing methods, however, are constrained to identifying models within a static set included in the classifier training, failing to adapt to newly emerged unseen models dynamically. To bridge this gap, we aim to develop a generalized model fingerprint extractor capable of zero-shot attribution, effectively attributes unseen models without exposure during training. Central to our method is a model synthesis technique, which generates numerous synthetic models mimicking the fingerprint patterns of real-world generative models. The design of the synthesis technique is motivated by observations on how the basic generative model’s architecture building blocks and parameters influence fingerprint patterns, and it is validated through two designed metrics that examine synthetic models’ fidelity and diversity. Our experiments demonstrate that this fingerprint extractor, trained solely on synthetic models, achieves impressive zero-shot generalization on a wide range of real-world generative models, improving model identification and verification accuracy on unseen models by over 40% and 15%, respectively, compared to existing approaches.
@@ I Introduction
Figure 1: Illustration of existing model attribution method (above) and our method (below). Existing methods rely on training with a limited set of real-world models, allows them to attribute seen classical models included in the training data. However, they struggle with generalizing to emerging unseen models like Stable Diffusion and DALL-E due to a gap in fingerprint distribution. Our method, in contrast, trains on numerous synthetic models that replicate a broader range of real-world generative model fingerprints, enabling it to attribute unseen models effectively in a zero-shot manner.
@@
Existing research seeks to identify the unique fingerprints on the images they generate, which can be leveraged to attribute a generated image to its source model. A commonly adopted model attribution paradigm frames the task as a multi-class classification problem [6, 8, 9, 10]. In this setup, images generated by a limited and static set of models are used to train a classifier, where each image is labeled with a unique model ID. During testing, the test image comes from a seen model predefined in the training set, and the classifier would identify the model ID of the image based on the model fingerprint it has learned [6, 8, 9]. However, there arises a scenario where a testing image originates from an unseen model not present during training. To tackle this, some methodologies [10, 11] adopt an open-set setup, assigning them an “unknown” label.
Despite these efforts, conventional methods, trained on a static set of seen models, struggle to adapt to the newly emerged unseen models dynamically. This limitation is exemplified in Figure 1(a), which presents the t-SNE feature extraction results from an existing method [10]. We train this method on classical models such as GAN, VAE, and Flow models  [^1] Include all the GAN, VAE, and Flow models in Table I, and test on both these and the emerging Diffusion models. The results reveal a significant generalization gap when the method encounters new types of generative models. This gap stems from the restricted scope of seen models that can be practically collected, whose fingerprint distributions may significantly differ from those of newly emerging models. As generative technology continues to evolve, the variety of unseen models expands, further challenging the capacity of existing methods to adapt dynamically.
To bridge this gap, our goal is to develop a more generalized model fingerprint extractor capable of attributing unseen models efficiently in a zero-shot setting. Central to our approach is a model synthesis technique that creates plenty of synthetic models that mimic a broad range of fingerprint patterns in real-world generative models. To gain insights for replicating these patterns, our initial step is to explore why generative models display unique fingerprint patterns, which, although typically imperceptible in the spatial domain, manifest as distinct spectral patterns in high-frequency components. Through analyzing the architecture and parameters of generative models, we observe: 1) The type of basic network components such as upsampling, activation function, normalization, and parameters of convolution layers are key factors influencing the spectral patterns of the generated images. 2) Across the generative blocks within a large generative model, due to the upsampling layer that naturally attenuates earlier layers’ high-frequency components, the last few generative blocks are more influential in determining the output’s spectrum patterns.
We incorporate these observations into our model synthesis technique. Our approach is based on a shallow auto-encoder architecture, comprising fewer generative blocks compared to typical generative models. According to the discussion above, employing a synthesis architecture with a small number of generative blocks could ensure fidelity in mimicking fingerprints. We then increase the architecture diversity by varying the types of upsampling layers, activation functions, normalization layers, and the number and sequence of these layers. Additionally, we enhance the parameter diversity by altering training seeds. Consequently, this approach allows us to get 5760 synthetic models across 288 different architectures by minimizing the reconstruction loss. Owing to the simple architecture and training objective, training these synthetic models requires much less time than real generative models. Subsequently, we leverage these synthetic models to train the fingerprint extractor, to extract fingerprints from these models, and distinguish between each other. We combine classification and metric loss to enhance the discrimination of learned fingerprint embedding.
We establish two metrics to verify the effectiveness of our model synthesis strategy in terms of fidelity and diversity. The former assesses the authenticity of fingerprint pattern mimicking by comparing the spectrum distribution between synthetic and real models, particularly in high-frequency components. The latter evaluates the diversity of spectrum patterns of synthetic models by examining the extent to which the synthetic models occupy the spectrum representation space. The results indicate that the synthesis options we designed indeed enhance the distribution fidelity and diversity of spectrum patterns.
@@
Experimental results demonstrate that our fingerprint extractor, despite being solely trained on synthetic models, exhibits strong zero-shot attribution capabilities on a broad scope generative models, including the classical GAN, VAE, Flow models, and emerging Diffusion models like Stable Diffusion and DalleE-3. Partial visualization results are shown in Figure 1(b). In the model attribution scenarios, including model identification and verification on unseen models, our method significantly outperforms existing approaches, achieving accuracy improvements of over 40% and 15% respectively.
@@
1) We propose to address a practical problem in applying model attribution in the real world, to expand the attribution target to unseen models in a zero-shot setting.
2) We propose to solve the zero-shot model attribution problem based on training on numerous synthetic models, which mimic the fingerprint patterns of real-world generative models. The synthesis strategy is motivated by observations on how generative model architecture building blocks and parameters influence fingerprint patterns and is validated through two designed metrics that examine synthetic models’ fidelity and diversity.
@@ II-A Model Attribution
In essence, model attribution methods rely on some relationship between the internal states of generative models and the generated images. Depending on how such a relationship is created and identified, existing studies could be classified into three types: watermark-based methods, model-inversion-based methods, and fingerprint-based methods.
Watermarking has been widely used in copyright protection for digital contents [12, 13], database [14, 15], and recently, generative models [16, 17, 18, 19]. Watermarking-based model attribution methods include those based on training data embedding [16, 17], weight modulation [19, 18], and latent vector watermarking [19]. Although watermarks facilitate model attribution, the injected watermark can hurt the quality of the generated images, besides, the attribution target is limited to models that have been watermarked, rather than any model, which limits its flexibility and application scenarios. Model-inversion-based methods [20, 21, 22] typically involves reversing the process to map an image back to its original latent space representation, followed by an attempt to regenerate the image. This technique is applicable only to white-box models, as it relies on accessing the model’s parameters.
Fingerprint-based methods attribute models by identifying unique fingerprints on images of generative models, which are particularly evident in the spectrum. Compared to the two previously mentioned techniques, fingerprint-based model attribution approaches utilize the distinctive characteristics of generative models without requiring the injection of information into the model. They can be applied to both white-box and black-box models by relying solely on the output images. Marra et al. [5] used averaged noise residuals to represent model fingerprints and found that these fingerprints are periodic. Subsequent works [6, 8, 9] confirmed the existence of these fingerprints and achieved high accuracy with a fixed, finite set of models following a closed-set classification approach. Asnani et al. [23] extended this research to trace the architectural components of models that generate images, which is beyond the scope of this paper. In open-world scenarios, with an unlimited number of unknown models, recent works [24, 10, 11] addressed model attribution using an open-set approach. However, the challenge of generalized model attribution remains unresolved. Current solutions are hampered by the seen/unseen imbalance issue in real-world scenarios, where only a limited number of seen models can be sampled while the number of unseen models continues to grow. Finding ways to improve the generalization performance of model attribution to new, emerging models remains a challenge.
@@ II-C Pre-train on Synthetic Data
SNGAN
ProGAN
Figure 3: Spectral effects of different types of upsampling layers.
III Method
In this work, our primary goal is to design a fingerprint extractor that could generalize to unseen models in the open world. To accomplish this goal, we propose a new approach by utilizing synthetic models, to mimic the fingerprint patterns of real-world generative models. By this approach, we significantly broaden the scope of the training data of fingerprint extractor, and consequently diminish the generalization gap. In the following, we begin with an analysis of the factors influencing the fingerprint patterns of generative models, detailed in Section III-A. Based the analysis, we then design our model synthesis strategy in Section 8. Finally, in Section III-C, we use the synthetic models to train the fingerprint extractor to perform the model attribution task.
III-A Preliminary Analysis of Model Fingerprint
@@
Our analysis mainly focuses on the generator (or decoder) part of CNN-based generative models. These generators produce images from either an encoded or a predefined latent vector. Typically, a generator comprises multiple generative blocks that transform the input latent vector into images. Common operations within the generative block involve components such as upsampling, convolution, normalization, and activation function. In the following, we aim to discuss the impact of these basic components and different generative blocks on the spectrum patterns of output images. Before the analysis, we first provide the background about analysis tools that will be employed in this section.
Background. Our analysis is based on the Discrete Fourier Transform (DFT). Specifically, for the input 2D image $I\in\mathbb{R}^{M\times N}$, the discrete Fourier transform for $I$ is denoted by:
$F(u,v)=\sum_{x=0}^{M-1}\sum_{y=0}^{N-1}I(x,y)\cdot e^{-i2\pi(ux/M+vy/N)},$
@@
where $F(u,v)$ denotes the coefficient under the frequency component $u$ and $v$, along the horizontal and vertical axis of the image. These coefficients consist of the so-called Fourier spectrum.
Sometimes, it is more convenient to consider the reduced 1D power spectrum [26], which is a simple yet characteristic 1D representation of the 2D Fourier spectrum. Technically, it is got by Azimuthal Integration (AI) over radial frequencies on the 2D Fourier spectrum:
$AI(\omega)=\int_{0}^{2\pi}\left\|F\left(\omega\cdot\cos(\phi),\omega\cdot\sin(\phi)\right)\right\|^{2}\mathrm{~{}d}\phi,$ (1)
where each value of $AI(\omega)$ represents the radial integral of the power over the 2D spectrum at a given radius $\omega$. Consequently, the reduced 1D power spectrum effectively captures the distribution of different frequency components without regard to angular direction.
SNGAN
ProGAN
Figure 4: Spectral effects of different types of activation functions.
Network Components. We first study the architecture influences, including the types of three key architecture components: upsampling, activation function, and normalization. Subsequently, we discuss the parameter influences associated with convolution layers.
  • Upsampling Layers. Upsampling layers serves to gradually increase the resolution of a generated image. Common types of upsampling layers are interpolation-based upsampler such as nearest and bilinear, and transposed convolution. Different types of upsampling are discussed in existing works to leave different spectral properties [25, 27]. To demonstrate this, we trained ProGAN [35] and SNGAN [36] models on the CelebA dataset, each with three types of upsampling layers. Figure 3 presents the reduced 1D power spectrum of these models, where the x-axis indicates different frequencies from low to high, and the y-axis denotes the power of these frequency components. The figure illustrates that nearest and bilinear, which act as low-pass filters, significantly reduce high-frequency content, resulting in lower power at higher frequencies on the spectrum’s right side. In contrast, transposed convolution, which can be conceptualized as zero-interleaving followed by a convolution with a learnable kernel, tends to preserve more high-frequency components. This is because the learned convolution kernels in transposed convolution layers are not guaranteed to be low-pass, allowing more high-frequency artifacts to remain in the generated images [25].
  • Activation Function. The choice of activation functions introduces non-linearities that shape the model’s frequency response. Different functions may enhance or suppress different aspects of the frequency spectrum. To verify this, we trained ProGAN and SNGAN models on the CelebA dataset, each with three different activation functions. As shown in Figure 4222To make the high-frequency discrepancy more evident, we first use a denoising filter by [37] with noise parameter $\sigma=1$ to remove the low-frequency semantic contents., generative models with ReLU activation generate images with more high-frequency components than Sigmoid and Tanh, as indicated by higher values on the right side of the reduced spectrum. The results may because that Tanh and Sigmoid compress the output into a bounded range, which has smooth properties and can damping high-frequency components. While ReLU is linear for all positive values, and zero otherwise, which allows high-frequency components to pass through unaffected in positive regions.
    Figure 5: Spectral effects of different types of normalization functions.
  • Normalization. Normalization techniques adjust data within the network by normalizing based on calculated means and variances, which are computed differently depending on the method, such as across the entire batch for Batch Normalization [38] or for each individual sample in the case of Instance Normalization [39]. Despite employing similar mathematical formulations, these techniques could influence the network’s behavior differently. Due to the variations in the optimization landscapes they create, different types of normalization may lead to distinctive distributions of parameters, and consequently different spectrum patterns. As demonstrated in Figure 5, we train six SNGAN models using two types of normalization, each type with three different initial seeds. As seen, models with the same normalization type often exhibit more consistent spectral patterns compared to those trained using different normalization methods. This underscores the impact of normalization choices on the diversity of spectrum patterns of generative models.
Convolution Parameters. Many factors, including optimization, initialization, and other configurations, can influence the parameters of a model. Our study primarily focuses on the parameters of convolution layers, which constitute a significant portion of a CNN-based model’s parameters. The shapes and sizes of convolution kernels critically determine how frequency components are handled. Different kernels can lead to different emphases on certain frequency bands, thus altering the model’s overall spectrum pattern. This phenomenon is illustrated in Figure 6, which shows the spectrum of two StyleGAN2 [40] models that have identical architecture but differ in their parameters, trained respectively on the FFHQ and FFHQu datasets. For this analysis, we extracted the parameters from a specific output channel in the last convolutional layer of each model’s final generative block. We then plot two figures for each model: the summed spectrum of the convolution kernels (left), and the spectrum of the output feature map for this channel (right). As demonstrated in the figure, for both models, the spectrum of the output feature map closely mirrors the spectrum of the convolution parameters. This alignment highlights how the convolutional layer’s kernel characteristics directly shape the spectrum pattern of the output.
(a) StyleGAN2 (FFHQ)
(b) StyleGAN2 (FFHQu)
Figure 6: Spectral effects of convolution parameters. This figure compares the summed spectrum of the convolution parameters with the spectrum of the output feature map.

### v3

@@ Abstract
Nowadays, generative models are shaping various fields such as art, design, and human-computer interaction, yet they are accompanied by copyright infringement and content management challenges. In response, existing research seeks to identify the unique fingerprints on the images they generate, which can be leveraged to attribute the generated images to their source models. However, existing methods are restricted to identifying models within a static set included in classifier training, incapable of adapting dynamically to newly emerging unseen models. To bridge this gap, this paper aims to develop a generalized model fingerprint extractor capable of zero-shot attribution that effectively attributes unseen models without exposure during training. Central to our method is a model synthesis technique, which generates numerous synthetic models that mimic the fingerprint patterns of real-world generative models. The design of the synthesis technique is motivated by observations on how the basic generative model’s architecture building blocks and parameters influence fingerprint patterns, and it is validated through designed metrics to examine synthetic models’ fidelity. Our experiments demonstrate that the fingerprint extractor, trained solely on synthetic models, achieves impressive zero-shot generalization on a wide range of real-world generative models, improving model identification and verification accuracy on unseen models by over 40% and 15%, respectively, compared to existing approaches.
@@ I Introduction
Figure 1: Illustration of the existing methods (above) and ours (below). Existing methods depend on training with a limited set of real-world models, struggling to distinguish emerging unseen models like Stable Diffusion (SD) and DALL-E series. In contrast, our method trains on numerous synthetic models that replicate a broader range model fingerprints, enabling it to distinguish and attribute unseen models effectively.
@@
Model attribution methods identify unique fingerprints in images generated by source models. These methods can leverage these fingerprint features to trace a generated image back to its originating model. A commonly adopted model attribution paradigm frames the task as a multiclass classification problem [6, 8, 9, 10]. In this setup, images generated by a limited and static set of models are used to train a classifier, where each image is labeled with a unique model ID. During testing, a test image comes from a seen model in the training set, and the classifier would identify the model ID based on the model fingerprint it has learned [6, 8, 9]. This framework looks promising, but has a major limitation: classical methods fail to attribute models not present in the training data, assigning incorrect classification labels to such new models. This drawback is unacceptable in today’s rapidly evolving landscape, where new models emerge quickly. Although some works [10, 11] consider an open-set setup by assigning an “unknown” label for new models, these approaches still struggle to adapt to the newly emerged unseen models dynamically.
To tackle the flexibility challenge, we aim to develop a more generalized model attribution method capable of zero-shot model attribution, efficiently attributing unseen models without requiring any training on samples from these models. Instead of relying on a classification head to classify the generated image to a static set of seen models that have been included in training, we view model attribution as comparing the ”fingerprint” distance between the test sample and each model’s fingerprint within a model gallery. This framework allows us to easily deal with newly emerged models, we could simply collect samples from new models, use the fingerprint extractor to create their fingerprints, and dynamically updates to the model gallery, without requiring any training on these models.
How to train an effective fingerprint extractor capable of generalizing on unseen models? Traditional wisdom suggests training it on a limited set of real-world models. However, this approach suffers from poor generalization on unseen models, which is exemplified in Figure 1 (above). To overcome this, we propose pre-training the extractor with a significantly larger set of synthetic models (e.g., 100 times more in our experiments), which provides similar fingerprint patterns as real-world models but much more diverse.
Our model synthesis strategy is based on analysis in the frequency domain, which is inspired by previous works [12, 13] demonstrating that fingerprints are unobservable in the spatial domain but prominent in the frequency domain. To understand how unique fingerprint patterns of real-world generative models are formed, we examine the mechanism from the perspective of basic building blocks of generative models. Our findings reveal that:
  • The type of basic network components—such as upsampling, activation functions, normalization, and convolution layer parameters—significantly influences the spectral patterns of generated images.
  • Due to the upsampling layer naturally attenuating high-frequency components from earlier layers, the last few generative blocks within a large model are more influential in determining the output’s spectral patterns.
Based on these insights, we hypothesize that synthesizing shallow generative blocks with various combinations of basic components can effectively represent the fingerprint space of unseen models. Therefore, we develop a simple yet effective model synthesise strategy.
Inspired by the above analysis, we construct the synthetic models based on a shallow auto-encoder architecture, comprising fewer generative blocks compared to typical generative models. We increase synthesis diversity by varying the basis components, including the types of upsampling layers, activation functions, normalization layers, the number and sequence of these layers, and training seeds. Consequently, we get 5760 synthetic models across 288 different architectures by minimizing the reconstruction loss. This model synthesis strategy has advantages in terms of fidelity and efficiency. By comparing the spectrum distribution between synthetic and real-world models, our results indicate that our synthesis options allow the spectrum distribution of synthetic models to more closely align with those of real-world generative models. Further, synthesizing a model takes only 83 seconds on average with a single 3090 GPU. This is remarkable faster than the naive solution of model synthesis (i.e., training state-of-the-art generative models from scratch). For example, it takes 14 days 22 hours to train a StyleGAN [14] model that generate with a 256$\times$256 resolution [^1] https://github.com/NVlabs/stylegan, which is significantly slower than ours.
@@
We leverage these synthetic models to train the fingerprint extractor, to extract fingerprints from these models, and distinguish between each other. We combine classification and metric loss to enhance the discrimination of learned fingerprint embedding. Experimental results demonstrate that: 1) Our fingerprint extractor, despite being solely trained on synthetic models, exhibits strong zero-shot attribution capabilities on a broad scope generative models, including the classical GAN, VAE, Flow models, and emerging Diffusion models like Stable Diffusion and DalleE-3. Partial visualization results are shown in Figure 1 (below). 2) We consider two model attribution scenarios, including model identification and model verification. Our method significantly outperforms existing approaches on unseen models, achieving accuracy improvements of over 40% and 15% respectively. 3) Our method is capable of tracing the LoRA-variants back to their base models only relying on the generated images, providing an efficient defense tool against model intellectual property infringements.
@@
1) We propose to solve the problem of zero-shot model attribution, which expands the model attribution target to unseen models that are not included in training.
2) We propose solving the zero-shot model attribution problem by training on numerous synthetic models that mimic the fingerprint patterns of real-world generative models. Our synthesis strategy is inspired by observations on how generative model architecture building blocks and parameters influence fingerprint patterns. This strategy has advantages in terms of fidelity and efficiency.
@@ II-A Model Attribution
In essence, model attribution methods rely on some relationship between the internal states of generative models and the generated images. This paper mainly focuses on fingerprint-based model attribution, which relies on the unique fingerprint patterns on the output images of generative models. Marra et al. [5] used averaged noise residuals to represent model fingerprints and found that these fingerprints are periodic. Subsequent works [6, 8, 9] confirmed the existence of these fingerprints and achieved high accuracy with a fixed, finite set of models following a closed-set classification approach. Asnani et al. [15] extended this research to trace the architectural components of models that generate images, which is beyond the scope of this paper. In real-world scenarios, images often originate from unseen models not present during training. Recent works [10, 11] addressed this by using an open-set approach, which attributes seen models and rejects unseen models not included in training. However, as generative technology continues to evolve, the variety of unseen models expands continuously, challenging the capacity of existing methods to adapt in the real-world dynamically. To bridge this gap, our goal is to develop a more generalized model fingerprint extractor capable of zero-shot model attribution, efficiently attributing unseen models without requiring any training on samples from these models. Table I summarizes the differences between our method and related works mentioned above.
TABLE I: Differences between our method and related works. Our method enables zero-shot model attribution, allowing us to attribute unseen models without needing training samples from them. In contrast, other methods can only attribute seen models that were included in training.
Method Task Training Data Type Number of Training Models Goal
[5, 6, 8, 9] Closed-set attribution Real-world models $<$ 50 Attribute seen models
[10, 11] Open-set attribution Real-world models $<$ 50 Attribute seen models, reject unseen models
Ours Zero-shot attribution Synthetic models $>$ 5000 Attribute unseen models
@@ II-C Pre-train on Synthetic Data
III Method
In this work, our primary goal is to design a fingerprint extractor that could generalize to unseen models in the open world. To accomplish this goal, we propose a new approach by utilizing synthetic models, to mimic the fingerprint patterns of real-world generative models. By this approach, we significantly broaden the scope of the training data of fingerprint extractor, and consequently diminish the generalization gap. In the following, we begin with an analysis of the factors influencing the fingerprint patterns of generative models, detailed in Section III-A. Based the analysis, we then design our model synthesis strategy in Section 4. Finally, in Section III-D, we use the synthetic models to train the fingerprint extractor to perform the model attribution task.
III-A Preliminary Analysis of Model Fingerprint
@@
The common architecture of an image generative model typically includes several key components designed to generate high-quality images from a latent representation or input data. Some of the most widely used architectures in image generative models include Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), Diffusion Models, etc. Image generative models, regardless of their specific type, typically consist of a series of generative blocks. Each block is designed to progressively transform the noise/latent vector into an image through several operations. To provide a cohesive mathematical foundation, consider a common generative block architecture consisting of upsampling, convolution, activation, and normalization layers applied to a Gaussian noise input $\mathbf{z}$. The block’s output $\mathbf{y}$ can be expressed as:
$\mathbf{y}=\mathcal{N}\left(\phi\left(F\left(U\mathbf{z}\right)\right)\right),$ (1)
where:
  • $U$: Upsampling operator.
  • $F$: Convolution operator with kernel $\mathbf{W}_{conv}$ and bias $\mathbf{b}$.
  • $\phi$: Nonlinear activation function.
  • $\mathcal{N}$: Normalization function.
In the following, we aim to discuss the impact of these basic components and different generative blocks on output’s spectrum patterns. To validate the theoretical insights, we also conducted experiments using ProGAN, SNGAN, and StyleGAN2 models with varying network components in Appendix -A, analyzing their spectrum influence on output images.
III-A1 Upsampling Effects
In the spatial domain, the upsampling operation $U\textbf{z}$ is equivalant to zero-interleaving the input signal z and then applying a standard convolutional operation, which can be expressed as:
$U\textbf{z}=\text{ZeroInter}(\textbf{z})\otimes\textbf{K}_{\text{up}},$ (2)
@@ III-A1 Upsampling Effects
Using the Convolution Theorem, the upsampling operation in the frequency domain becomes:
$\mathcal{F}\{U\textbf{z}\}(\omega)=\text{Repeat}_{2,2}(\mathcal{F}\{\textbf{z}\})(\omega)\cdot\mathcal{F}\{\textbf{K}_{\text{up}}\}(\omega),$ (3)
where $\text{Repeat}_{2,2}$ represents the replication of the input spectrum $\mathcal{F}\{\textbf{z}\}$ along two frequency dimensions by a factor of 2. $\mathcal{F}\{\textbf{K}_{\text{up}}\}$ represents the Fourier Transform of the upsampling kernel $\textbf{K}_{\text{up}}$. $\cdot$ denotes pointwise multiplication in the frequency domain.
Fingerprint: The upsampling kernel $\textbf{K}_{\text{up}}$ determines how the replicated spectra are weighted. Its magnitude |$|\mathcal{F}\{\textbf{K}_{\text{up}}\}(\omega)|$ directly shapes the output signal. Different types of upsampling are discussed in existing works to leave different spectral properties [16, 18]. Basic kernels like nearest-neighbor or bilinear correspond to low-pass frequency responses, attenuating high-frequency content. In contrast, learned transposed convolution kernels can either preserve, boost, or selectively filter out certain frequency bands. Thus, they tends to preserve more high-frequency components [16]. Thus, the upsampling ”fingerprint” represents a characteristic spectral pattern formed through frequency shaping by the upsampling kernel.
III-A2 Convolution Effects
Applying the Convolution Theorem, convolution $F$ in the spatial domain translates to multiplication in the frequency domain:
$\mathcal{F}\{F\mathbf{x}\}(\omega)=\mathcal{F}\{\mathbf{W}_{conv}\}(\omega)\cdot\mathcal{F}\{\mathbf{x}\}(\omega)+\mathbf{b}\cdot\delta(\omega),$ (4)
where $\mathbf{x}$ is the input signal, $\mathbf{W}_{conv}$ and $\mathbf{b}$ are the kernel and bias of the Convolution operator.
Fingerprint: The output spectrum is shaped by |$|\mathcal{F}\{\mathbf{W}_{conv}\}(\omega)|$. Each learned convolution kernel acts as a frequency-selective filter, enhancing certain frequency bands and suppressing others. This creates a spectral “fingerprint” dictated by the learned convolution kernel’s frequency response.
III-A3 Nonlinear Activation Effects
In the spatial domain, applying $\phi(\mathbf{x})$ is elementwise. While in frequency domain, elementwise nonlinearities do not translate to simple pointwise operations. Instead, nonlinearities correspond to convolution-like mixing of frequencies. To simplify the analysis, the nonlinearity can be approximated by a Taylor series expansion. For small values of $\mathbf{x}$, common nonlinear activation functions like ReLU, Sigmoid and Tanh can be approximated using their Taylor series expansions:
  • ReLU: The expansion is [^2] Since ReLU is non-differentiable at $\mathbf{x}=0$, we approximate it using $\text{softplus}_{10}(x)$, and use the Taylor series expansions of $\text{softplus}_{10}(x)$ for ReLU.:
    $ReLU(\mathbf{x})\approx\frac{\log(2)}{10}+\frac{1}{2}\mathbf{x}-\frac{5}{4}\mathbf{x}^{2}+\frac{5}{48}\mathbf{x}^{3}+\mathcal{O}(\mathbf{x}^{4}).$ (5)
  • Tanh: The expansion is:
    $\tanh(\mathbf{x})\approx\mathbf{x}-\frac{\mathbf{x}^{3}}{3}+\mathcal{O}(\mathbf{x}^{5}).$ (6)
  • Sigmoid: The expansion is:
    $\text{sigmoid}(\mathbf{x})\approx\frac{1}{2}+\frac{\mathbf{x}}{4}-\frac{\mathbf{x}^{3}}{48}+\mathcal{O}(\mathbf{x}^{5}).$ (7)
In the frequency domain, higher-order terms (e.g., $\mathbf{x}^{2}$) result in repeated convolutions of the input signal’s Fourier Transform $H(\omega)$ with itself:
$\mathcal{F}\{\mathbf{x}^{n}\}(\omega)=\underbrace{H(\omega)*H(\omega)*\cdots*H(\omega)}_{n\text{ times}},$ (8)
where $H(\omega)=\mathcal{F}\{\mathbf{x}\}(\omega)$. This convolution operation generates harmonics at integer multiples of the input frequency $\omega$. For example, $H(\omega)*H(\omega)$ introduces even-order harmonics, such as $2\omega$, and a DC component ($\omega$=0).
Fingerprint: The frequency-domain “fingerprint” of the nonlinear activation is the introduction of harmonics and frequency mixing patterns, transforming a input into a broader spectral distribution. For instance, ReLU introduces strong even harmonics due to the quadratic term ($\frac{5}{4}\mathbf{x}^{2}$). Tanh and Sigmoid add weak odd harmonics ($\frac{\mathbf{x}^{3}}{3}$ and $\frac{\mathbf{x}^{3}}{48}$).
III-A4 Normalization Effects
Normalization techniques adjust data by normalizing its mean and variance. In the frequency domain, different types of normalization have the same formula as below:
$\mathcal{F}\{\hat{\mathbf{x}}\}(\omega)=\frac{\gamma}{\sigma}\mathcal{F}\{\mathbf{x}-\mu\}(\omega)+\mathcal{F}\{\beta\}(\omega),$ (9)
where $\mu$ and $\sigma$ are the mean and standard deviation. $\gamma$ and $\beta$ are learnable scale and shift parameters. $\mu$ and $\sigma$ are computed differently depending on the method, such as across the entire batch for Batch Normalization [25] or for each individual sample in the case of Instance Normalization [26].
Fingerprint: Despite having the same frequency formulation, due to the variations in the optimization landscapes they create, different types of normalization can lead to distinctive distributions of parameters [26], and consequently different spectrum patterns. [^3] For instance, Figure 14 in Appendix -A shows that models trained with different seeds but the same normalization type generally exhibit more consistent spectral patterns than those trained with a different normalization method.